OpenStudy (anonymous):

The data in the table listed below relates the distance a ball travels down a ramp to the time that has passed since it was released. Find a variation equation that fits the data. Time, t(sec)| 0,1,2,3,4,5 Distance, d(in)| 0, 0.4, 1.6, 3.6, 6.4, 10.0

OpenStudy (aum):

The x-values (Time) are increasing by 1. Take the differences in the y-values (Distance): 0.4, 1.2, 2.0, 2.8, 3.6 Since the y-values are NOT increasing by the same amount each time the variation is NOT linear. Take the differences one more time: 0.8, 0.8, 0.8, 0.8. The second differences are constant and therefore this is a quadratic variation.

OpenStudy (anonymous):

Soooo could you give me an example of what you would write for an variation equation?

OpenStudy (anonymous):

@aum

OpenStudy (aum):

Try y = ax^2 when x = 1, y = a = 0.4 so a = 0.4 y = 0.4x^2 x = 1, y = 0.4 x = 2, y = 0.4 * 4 = 1.6 x = 3, y = 0.4 * 9 = 3.6 ....

OpenStudy (anonymous):

okay. that makes sense. so whats the equation you would create to describe this?

OpenStudy (aum):

Distance d = 0.4 * t^2

OpenStudy (aum):

You put t = 0, 1, 2, 3, 4,5 into the equation d = 0.4 * t^2 and you will get the d values given in the table.

OpenStudy (anonymous):

gotcha so your answer is d= 0.4*t^d"?

OpenStudy (aum):

\[\Large d = 0.4t^2\]